OpenStudy (anonymous):

the equations that are parallel and perpendicular to y = 2x - 1 and that pass through the point (4, 1)

OpenStudy (anonymous):

perpendicular means negative inverse of slope, since initial slope is 2, the slope of the perpendicular is -1/2 so the equation you are generating will be y=-1/2*x+b 1=-1/2*4+b 1=-2+b b=3 y=-1/2*x+3

OpenStudy (anonymous):

the parallel has the same slope, so y=2*x+b 1=2*4+b 1=8+b -7=b y=2*x-7

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sorry for the delay on that second one, didnt see you wanted both

OpenStudy (anonymous):

its okay i kinda figured the second one out but thank you any ways.